the stiffness 'S' of a rectangular beam varies jointly as the width and the cube of the depth. if the stiffness is 90 when the width is 2 and the depth is 3, find the stiffness for a beam with a width of 5 and a depth of 2
Answer by Alan3354(69443)

the stiffness 'S' of a rectangular beam varies jointly as the width and the cube of the depth. if the stiffness is 90 when the width is 2 and the depth is 3, find the stiffness for a beam with a width of 5 and a depth of 2
-----------------
S = k*w*d^3 (k is a constant)
90 = k*2*27
k = 90/54 = 5/3
---------------
S = (5/3)*5*2^3
S = 200/3